OpenStudy (anonymous):

I WILL MEDAL AND FAN! What is the discriminant D of the equation? x^2 – 2x + 2 = 0 D = _________

OpenStudy (faiqraees):

For the general equation of quadratic equation \[\large\rm ax^2+bx+c=0 \]the discriminant is \[\large\rm d=\sqrt{b^2-4(a)(c)} \]

OpenStudy (anonymous):

so the answer is 4? @FaiqRaees

OpenStudy (superteenblue):

@Kyliejones

OpenStudy (alivejeremy):

It -4

OpenStudy (alivejeremy):

The answer

OpenStudy (anonymous):

okay thanks!!

OpenStudy (alivejeremy):

Yw.......Anytime

OpenStudy (faiqraees):

no the answer is √-4

OpenStudy (alivejeremy):

Still -4

OpenStudy (alivejeremy):

-_-

OpenStudy (anonymous):

thanks guys!!

OpenStudy (alivejeremy):

Yw again

OpenStudy (faiqraees):

oh sorry there should be no under root sign it is just -4

OpenStudy (faiqraees):

\[\large\rm d=b^2-4(a)(c)\]